\newpage
\section{References}
\begin{thebibliography}{99}

%CCCCCCCCCCCCCCCCCCCCCCCCCCC
%\bibitem{CC2009} Cobeli Cristian. \textit{\color{blue}On a Problem of Mordell with Primitive Roots}. http://arxiv.org/abs/0911.2832.


%\bibitem{CZ1998} Cobeli, Cristian; Zaharescu, Alexandru. \textit{\color{blue}On the distribution of primitive roots mod $p$}. Acta Arith. 83 (1998), no. 2, 143-153.
%MR1490645.

\bibitem{CN2017} Carella, N. \textit{\color{blue}Densities Of Primes And Primitive Roots}. http://arxiv.org/abs/1707.06517.

\bibitem{CP2005} Crandall, Richard; Pomerance, Carl. \textit{\color{blue}Prime numbers. A computational perspective}. Second edition. Springer, New York, 2005.

\bibitem{CR1910} Carmichael, R. D. \textit{\color{blue}Note on a new number theory function}. Bull. Amer. Math. Soc. 16 (1910), no. 5, 232-238.

\bibitem{CZ1998} Cobeli, Cristian; Zaharescu, Alexandru. \textit{\color{blue}On the distribution of primitive roots mod $p$}. Acta Arith. 83, (1998), no. 2, 143-153.
%DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD
\bibitem{DH1937} Davenport, H. \textit{\color{blue}On Primitive Roots in Finite Fields}. Quarterly J. Math. 1937, 308-312. 

\bibitem{DS2012} Dietmann, Rainer; Elsholtz, Christian; Shparlinski, Igor E. \textit{\color{blue}On Gaps Between Primitive Roots in the Hamming Metric}. http://arxiv.org/abs/1207.0842. 
%EEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE
\bibitem{ES1957} Erdos, Paul; Shapiro, Harold N. \textit{\color{blue}On The Least Primitive Root Of A Prime}. 1957, euclidproject.org.
%FFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFF
%\bibitem{FS2001} Friedlander, John B.; Shparlinski, Igor E. \textit{\color{blue}Double exponential sums over thin sets}. Proc. Amer. Math. Soc. 129 (2001),  no. 6, 1617-1621.	
%MR1814088.
%\bibitem{FS2000} Friedlander, John B.; Hansen, Jan; Shparlinski, Igor E. \textit{\color{blue}Character sums with exponential functions.} Mathematika 47 
%(2000),  no. 1-2, 75-85 (2002).	MR1924489.

%GGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGG
%\bibitem{GZ2005} Garaev, M. Z. \textit{\color{blue}Double exponential sums related to Diffie-Hellman distributions}. Int. Math. Res. Not. 2005,  no. 17, 1005-1014. 
%MR2145707.
%\bibitem{GK2005} Garaev, M. Z.  Karatsuba, A. A. \textit{\color{blue}New estimates of double trigonometric sums with exponential functions.} http://arxiv.org/math/0504026. 	 
%MR2246404.

%HHHHHHHHHHHHHHHHHHHHHHHHHHHHH
\bibitem{HC1967} Hooley, C. \textit{\color{blue}On Artins conjecture.}  J. Reine Angew. Math. 225, 209-220, 1967.
%MR0207630.

%\bibitem{HG2007} Harman, Glyn. \textit{\color{blue}Prime-detecting sieves.} London Mathematical Society Monographs Series, 33. Princeton University Press, 
%Princeton, NJ, 2007. MR2331072
%\bibitem{HG1984}Hinz, Jurgen G. Some applications of sieve methods in algebraic number fields. Manuscripta Math.  48  (1984),  no. 1-3, 117-137. 

%IIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIII
%\bibitem{IK2004} Iwaniec, Henryk; Kowalski, Emmanuel. \textit{\color{blue}Analytic number theory.} AMS Colloquium Publications, 53. American Mathematical Society, Providence, RI, 2004.
%MR2061214.

%LLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLL
\bibitem{LE1878} Lucas, Edouard. \textit{\color{blue}Theorie des Fonctions Numeriques Simplement Periodiques}. (French) Amer. J. Math.  1  (1878),  no. 4, 289-321.
\bibitem{LN1997} Lidl, Rudolf; Niederreiter, Harald. \textit{\color{blue}Finite fields.} Second edition. Encyclopedia of Mathematics and its Applications, 20. Cambridge University Press, Cambridge, 1997.
%MR1429394.
%\bibitem{LH1977} Lenstra, H. W. \textit{\color{blue}On Artin’s conjecture and Euclid’s algorithm in global fields}, Invent. Math. 42 (1977), 201–224. MR0480413.


%KKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKK
\bibitem{KW1983} Keller, W. \textit{\color{blue}Factors of Fermat numbers and large primes of the form $k \dot 2n + 1$}. Math. Comp. 41 (1983), 661-673.



%MMMMMMMMMMMMMMMMMMMMMMMMMMMMMM
\bibitem{MO2021} MathOverflow. \textit{\color{blue}Least prime in Artin's primitive root conjecture}. http://mathoverflow.net/questions/386615.

\bibitem{MP2007} Moree, P. \textit{\color{blue}Artin prime producing quadratics}. Abh. Math. Sem. Univ. Hamburg  77  (2007), 109-127.
%\bibitem{ML1972} Mordell, L. J. \textit{\color{blue}On the exponential sum $\sum_{1\leq x \leq X} exp (2\pi i(ax+bg^x )/p)$}. Mathematika 19  (1972), 84-87.
%MR0318073.
%\bibitem{MJ2012} Maynard, James. \textit{\color{blue}On the Brun-Titchmarsh theorem}. Acta Arith. 157  (2013),  no. 3, 249-296.
%MR3019418.
%\bibitem{MP2004} Moree, Pieter. \textit{\color{blue}Artin's primitive root conjecture -a survey}. http://arxiv.org/math/0412262.
%MR3011564.
%\bibitem{MP1999} Moree, Pieter. \textit{\color{blue}On primes in arithmetic progression having a prescribed primitive root. } J. Number Theory 78  (1999),  no. 1, 85-98.	
%MR1706921.

%\bibitem{MV2007} Montgomery, Hugh L.; Vaughan, Robert C. \textit{\color{blue}Multiplicative number theory.} I. \textit{Classical theory}. Cambridge University Press, Cambridge, 2007.
%MR2378655.
%PPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP
\bibitem{PS1995}  Pappalardi, Francesco; Shparlinski, Igor. \textit{\color{blue}On Artin's conjecture over function fields}. Finite Fields Appl.  1  (1995),  no. 4, 399-404. 

%RRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRR
\bibitem{RH1994} Rose, H. E. \textit{\color{blue}A course in number theory}. Second edition. Oxford Science Publications. The Clarendon Press, Oxford University Press, New York, 1994.
%SSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSS
\bibitem{SP1969} Stephens, P. J. \textit{\color{blue}An average result for Artin conjecture.} Mathematika 16, (1969), 178-188.
http://doi.org/10.1112/S0025579300008159.
%\bibitem{SR1973} Stoneham, R. G. \textit{\color{blue}On the uniform e-distribution of residues within the periods of rational fractions with 
%applications to normal numbers}. Acta Arith. 22  (1973), 371-389. MR0318091.

%\bibitem{SP2003} Stevenhagen, Peter. \textit{\color{blue}The correction factor in Artin's primitive root conjecture}. Les XXII emes Journees Arithmetiques (Lille, 2001). J. Theor. Nombres Bordeaux 15 (2003), no. 1, 383-391.
%MR2019022.

%TTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTT
%\bibitem{TG2015} Tenenbaum, Gerald. \textit{\color{blue}Introduction to analytic and probabilistic number theory.} Translated from the Third French edition. American Mathematical Society, Rhode Island, 2015. MR3363366
%\bibitem{VR1973} Vaughan, R. C. \textit{\color{blue}Some applications of Montgomery's sieve}. J. Number Theory 5 (1973), 64-7.
%MR0342476.

%WWWWWWWWWWWWWWWWWWWWWWWWW
\bibitem{WJ1961} Wrench, John W. \textit{\color{blue}Evaluation of Artin's constant and the twin-prime constant}. Math. Comp. 15 1961 396-398.	
MR0124305.
\bibitem{WR2001} Winterhof, Arne. \textit{\color{blue}Character sums, primitive elements, and powers in finite fields}. J. Number Theory 91, 2001, no. 1, 153-163.
%\bibitem{ZD2009} Zywina, David. \textit{\color{blue}A refinement of Koblitz's conjecture}. http://arxiv.org/abs/0909.5280.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[top=3cm, bottom=3cm, left=3cm, right=3cm]{geometry}
\usepackage{tabu}
\usepackage{breqn}

\usepackage{amsmath,amsthm,amssymb,graphicx, multicol, array}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{hyperref}
\usepackage{setspace}
\usepackage{xcolor}
\usepackage{currfile}
\usepackage{mathrsfs}
\usepackage{multicol}

\usepackage[hang,flushmargin]{footmisc} 
\usepackage{tabto}

\DeclareMathOperator{\li}{li}
\DeclareMathOperator{\ord}{ord}

\DeclareMathOperator{\lcm}{lcm}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]

\newtheorem{conj}{Conjecture}[section]

\newtheorem{dfn}{Definition}[section]
\newtheorem{exe}{Exercise}[section]

\newcommand{\N}{\mathbb{N}}
\newcommand{\Z}{\mathbb{Z}}

\newcommand{\F}{\mathbb{F}}

\usepackage{fancyhdr}
\title{Prescribed Primitive Roots And The Least Primes}
\date{}
\author{N. A. Carella}

\begin{document}
\thispagestyle{empty}
\date{}

\maketitle
\textbf{\textit{Abstract}:} Let $q\ne \pm1,v^2$ be a fixed integer, and let $x\geq 1$ be a large number. The least prime number $p \geq3 $ such that $q$ is a primitive root modulo $p$ is conjectured to be $p\ll (\log q)(\log \log q)^3),$ where $\gcd(p,q)=1$. This note proves the existence of small primes $p\ll(\log x)^c$, where $c>0$ is a constant, a close approximation to the conjectured upper bound. \let\thefootnote\relax\footnote{ \today \date{} \\
\textit{AMS MSC2020}: Primary 11A07, 11N13, Secondary 11N05, 11N37.\\
\textit{Keywords}: Primitive root, Least prime number, Artin primitive root conjecture.}


\section{Introduction}\label{S3232.000}
It is a routine calculations, using the Laws of Quadratic Reciprocity, to verify that the subset of integers $\mathcal{Q}_0=\{q=5a+2\text{ or } q=5a+3:a\geq0\}$ are primitive roots modulo $p=5$. Here, $p=5$ is the least such prime. However, it is probably difficult or impossible to use the same analysis to verify which nonsquare integers in the subset of integers $\mathcal{Q}_1=\{q=5a+1\text{ or } q=5a+4:a\geq1\}$ are primitive roots of small primes $p\geq3$, which are bounded by a constant $p\leq c=c(q)$.

\begin{conj}   \label{conj3232.000}  {\normalfont (Granville, \cite{MO2021})} Let $q\ne 1, n^2$ be a fixed integer. Then, there exists a prime $p\ll (\log q)(\log \log q)^3$ such that $q$ is a primitive root modulo $p$.
\end{conj} 

In addition, the conjecture claims that for any prescribed nonsquared integer $q\geq2$, a small finite subset of Germain primes
\begin{equation} \label{eq3232.310}
\mathcal{G}=\{p=2r+1: r \text{ and } p \text{ are primes}\}=\{3, 5, 7,23,43,\ldots\}
\end{equation}
contains a least prime $p$ such that $q$ is a primitive root modulo $p$, with very  few exeptions, see \cite{MO2021}. The subset of Germain primes is suitable for this application because each of these primes has a very simple primitive root test. A weak version of this conjecture is considered in this note. The precise result is as follows.

\begin{theorem}   \label{thm3232.100}  Let $x \geq 1$ be a large number, and let $q\ne 1, n^2$ be a fixed integer. Then, there exists a prime $p\ll (\log x)^c$ such that $q$ is a primitive root modulo $p$, where $c>0$ is a constant.
\end{theorem} 

The preliminary notation, definitions, background results are discussed in Section \ref{S3242.100} to Section \ref{S3282.000}. Section \ref{S3292.000} presents a proof of Theorem \ref{thm3232.100}.

\section{Primitive Roots Tests}  \label{S3242.100}
For any prime $p \geq 3$, the multiplicative group $G$ of the prime finite fields $\mathbb{F}_p$ is a cyclic group of cardinality $p-1=\#G$. Similar result is true for any finite extension $\mathbb{F}_q$ of $\mathbb{F}_p$, where $q=p^k$ is a prime power. \\

\begin{dfn} \label{dfn3242.100} {\normalfont The \textit{multiplicative order} of an element $u\ne0$ in the cyclic group $\mathbb{F}_p^\times$ is defined by $\ord_p(u)=\min \{k \in \mathbb{N}: u^k \equiv 1 \bmod p \}$. An element is a \textit{primitive root} if and only if $\ord_p(u)=p-1$. 
}
\end{dfn}

The Euler totient function counts the number of relatively prime integers \(\varphi (n)=\#\{ k:\gcd (k,n)=1 \}\). This counting function is compactly expressed by
the analytic formula \(\varphi (n)=n\prod_{p \mid n}(1-1/p),n\in \mathbb{N} .\)\\

\begin{lem} {\normalfont (Fermat-Euler)} \label{lem3242.200}If \(a\in \mathbb{Z}\) is an integer such that \(\gcd (a,n)=1,\) then \(a^{\varphi (n)}\equiv
	1 \bmod n\).
\end{lem}

\begin{lem} \label{lem3242.500}  {\normalfont (Primitive root test in $\mathbb{F}_p$)} An integer $u \in \Z$ is a primitive root modulo an integer $n \in \N$ if and only if 
\begin{equation}\label{eq3242.500}
u^{\varphi(n)/p} -1\not \equiv 0 \mod  n
\end{equation}
for all prime divisors $p \mid \varphi(n)$.
\end{lem}
The primitive root test is a special case of the Lucas primality test, introduced in \cite[p.\ 302]{ LE1878}. A more recent version appears in \cite[Theorem 4.1.1]{CP2005}, and similar sources. \\

The Carmichael function is basically a refinement of the Euler totient function to the finite ring \(\mathbb{Z}/n \mathbb{Z}\). 

\begin{dfn} Given an integer
\(n=p_1^{v_1}p_2^{v_2}\cdots  p_t^{v_t}\), the Carmichael function is defined by
\begin{equation}
	\lambda (n)=\lcm \left (\lambda \left(p_1^{v_1}\right),\lambda \left (p_2^{v_2}\right ) \cdots  \lambda \left (p_t^{v_t}\right ) \right )
	=\prod _{p^v  \mid \mid  \lambda (n)} p^v,
\end{equation}
where the symbol \(p^v \mid \mid n,\nu \geq 0\), denotes the maximal prime power divisor of \(n\geq 1\), and 
\begin{equation}
	\lambda
	\left(p^v\right)= \left \{
	\begin{array}{ll}
		\begin{array}{ll}
			\varphi \left(p^v\right) 
			& \text{ if } p\geq 3\text{ or } v\leq 2,  \\
			2^{v-2} 
			& \text{ if } p=2 \text{ and }v \geq 3. \\
		\end{array}
		
	\end{array} \right . 
\end{equation}
\end{dfn} 
The two functions coincide, that is, \(\varphi(n)=\lambda (n)\) if \(n=2,4,p^m,\text{ or } 2p^m,m\geq 1\). And \(\varphi \left(2^m\right)=2\lambda
\left(2^m\right)\). In a few other cases, there are some simple relationships between \(\varphi (n) \text{ and } \lambda (n)\). In fact, it seamlessly
improves the Fermat-Euler Theorem: The improvement provides the least exponent \(\lambda (n) \mid  \varphi (n)\) such that \(a^{\lambda (n)}\equiv
1 \bmod n\). The ratio $\varphi (n)/\lambda (n) \geq1$ has many interesting properties studied in the literature.\\

\begin{lem} \label{lem2.2} { \normalfont (\cite{CR1910})}  Let \(n\in \mathbb{N}\) be any given integer. Then
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item The congruence \(a^{\lambda (n)}\equiv
1 \bmod n\) is satisfied by every integer \(a\geq 1\) relatively prime to \(n\), that is \(\gcd (a,n)=1\).

\item In every congruence \(x^{\lambda (n)}\equiv 1 \bmod n\), a solution \(x=u\) exists which is a primitive root \(\bmod  n\), and for any such
solution \(u\), there are \(\varphi (\lambda (n))\) primitive roots congruent to powers of \(u\).
\end{enumerate} 
\end{lem}

\begin{proof} (i) The number \(\lambda (n)\) is a multiple of every \(\lambda \left(p^v\right)=\varphi\left(p^v\right) \) such that \(p^v \mid  n\). Ergo, for any relatively prime integer \(a\geq 2\), the system of congruences 
\begin{equation}
a^{\lambda (n)}\equiv 1\bmod p_1^{v_1}, \quad a^{\lambda (n)}\equiv 1\bmod p_2^{v_2}, \quad \ldots, \quad a^{\lambda (n)}\equiv 1\bmod p_t^{v_t},
\end{equation}
where \(t=\omega (n)\) is the number of prime divisors in \(n\), is valid. 
\end{proof}

\begin{dfn}
	An integer \(u\in \mathbb{Z}\) is called a \textit{$\lambda$-primitive root} \(\text{mod } n \) if the least exponent \(\min  \left\{ m\in \mathbb{N}:u^m\equiv 1 \bmod n \right\}=\lambda
	(n)\).
\end{dfn}

\begin{lem} \label{lem3242.600}  {\normalfont (Primitive root test in $\mathbb{Z}/n\mathbb{Z}$)} An integer $u \in \Z$ is a primitive root modulo an integer $n \in \N$ if and only if 
\begin{equation}\label{eq3242.600}
u^{\lambda(n)/p} -1\not \equiv 0 \mod  n
\end{equation}
for all prime divisors $p \mid \lambda(n)$.
\end{lem}

\begin{lem} \label{lem3242.700}  {\normalfont (Primitive root lift)} Let $n$, and $u \in \mathbb{N}$ be integers, $\gcd(u,n)=1$. If $u$ is a primitive root modulo $p^k$ for each prime power divisor $p^k \mid n$, then, the integer $u \ne \pm 1, v^2$ is a primitive root modulo $n$.
\end{lem}

\begin{proof} Without loss in generality, let $n=pq$ with $p\geq 2$ and $q\geq 2$ primes. Let $u$ be a primitive root modulo $p$ and modulo $q$ respectively. Then
\begin{equation} \label{eq3242.700}
u^{(p-1)/r}  -1\not \equiv 0 \bmod  p  \qquad  \text{  and  } \qquad u^{(q-1)/s} -1 \not \equiv 0 \bmod  q,
\end{equation}
for every prime $r \mid p-1$, and every prime $s \mid q-1$ respectively, see Lemma \ref{lem3242.500}. Now, suppose that $u$ is not a primitive root modulo $n$. In particular,
\begin{equation}\label{eq3242.720}
u^{\lambda(n)/t} -1\equiv 0 \bmod  n  
\end{equation}
for some prime divisor $t \mid \lambda(n)$.\\

Let $v_t(\lambda(n))$, $v_t(p-1)$, and $v_t(q-1)$ be the $t$-adic valuations of these integers. Since $\lambda(n)=\lcm(\varphi(p-1), \varphi(q-1))$, it follows that at least one of the relations
\begin{equation}\label{eq3242.730}
v_t(\lambda(n))=v_t(p-1)  \qquad  \text{ or } \qquad v_t(\lambda(n))=v_t(q-1) 
\end{equation}
is valid. As consequence, at least one of the congruence equations
\begin{equation}\label{eq3242.740}
u^{\lambda(n)/t} -1\equiv 0 \bmod  n  \qquad  \Longleftrightarrow \qquad u^{\lambda(n)/t}-1 \equiv 0\mod p
\end{equation}
or
\begin{equation}\label{eq3242.750}
u^{\lambda(n)/t} -1\equiv 0 \bmod  n  \qquad  \Longleftrightarrow \qquad u^{\lambda(n)/t}-1 \equiv0 \mod q
\end{equation}
fails. But, this in turns, contradicts the relations in (\ref{eq3242.700}) that $u$ is a primitive root modulo both $p$ and $q$. Therefore, $u$ is a primitive root modulo $n$. 
\end{proof}

\begin{lem} \label{lem3242.850} The integer $2$ is a quadratic residue, {\normalfont (}quadratic nonresidue{\normalfont )} of the primes of the form $p=8k\pm1$, {\normalfont (}$p=8k\pm3$ respectively{\normalfont )}. Equivalently,
\begin{equation}\label{eq3242.850}
\left ( \frac{2}{p}\right )=(-1)^{\frac{p^2-1}{8}}.
\end{equation}
\end{lem}
\begin{proof}A detailed proof of the quadratic reciprocity laws appears in \cite[Theorem 1.5.]{RH1994}, and similar references.
\end{proof}
\begin{lem} \label{eq3242.860}  {\normalfont (Quadratic reciprocity law)} If $p$ and $q$ are odd primes, then
\begin{equation}\label{eq3242.870}
\left ( \frac{p}{q}\right )\left ( \frac{q}{p}\right )=(-1)^{\frac{p-1}{2}\frac{q-1}{2}}.
\end{equation}
\end{lem}
\begin{proof}A detailed proof of the quadratic reciprocity laws appears in \cite[Theorem 2.1.]{RH1994}, and similar references.
\end{proof}

\section{Very Short Primitive Roots Tests}  \label{S3252.200}
The set of Fermat primes 
\begin{equation}\label{lem3252.150}
\mathcal{F}=\left \{F_n=2^{2^n}+1:n\geq0\right \}= \left \{3,5,17,257,65537,\ldots\right \}
\end{equation}
has the simplest primitive root test: A quadratic nonresidue $q\geq3$ is a primitive root mod $F_n$. This follows from Lemma \ref{lem3242.500}. The next set of primes with a short primitive root test seems to be the set of generalized Germain primes.
\begin{dfn} \label{dfn3252.500} {\normalfont Let $s\geq 1$ be a parameter. The set of generalized Germain primes is defined by 
\begin{equation} \label{eq3252.340}
\mathcal{G}_s=\{p=2^s\cdot r+1: p \text{ and } r \text{ are primes}\}=\{3, 5, 7,13,17,23,29,37,43,41,73,\ldots\}\nonumber.
\end{equation}
}
\end{dfn}

\begin{lem} \label{lem3252.550} An integer $q\ne\pm1, n^2$ is a primitive root modulo a Germain prime $p=2^s\cdot r+1$ if and only if 
\begin{multicols}{2}
\begin{enumerate}[font=\normalfont, label=(\roman*)]
\item $\displaystyle  q^{2^{s-1}r}\not \equiv 1 \mod  p$,      
\item $\displaystyle q^{2^{s}}\not \equiv 1 \mod  p.$
\end{enumerate}
\end{multicols}
\end{lem}

\begin{proof} Let $p=2^sr+1$ be a Germain prime, where $r\geq 2$ is prime, and $s\geq1$ is an integer. Since the totient $p-1=2^sr$ has two prime divisors, an integer $q\ne\pm1, n^2$ is a primitive root modulo $p$ if and only if 
\begin{enumerate}[font=\normalfont, label=(\roman*)]
\item $\displaystyle  q^{(p-1)/2}=q^{2^{s-1}r}\not \equiv 1 \mod  p$,
\item $\displaystyle q^{(p-1)/r} =q^{2^{s}}\not \equiv 1 \mod  p,$
\end{enumerate}
see Lemma \ref{lem3242.500}.
\end{proof}

As the Germain primes, the set of primes of the form
\begin{equation}
\mathcal{A}=\{p=k\cdot 2^n+1:n\geq0\},    
\end{equation}
with $k\geq3$ a fixed prime, have a very short primitive root test, similar to the algorithm in Lemma \ref{lem3252.550}. Some of these primes are factors of Fermat numbers. There are many interesting problems associated with these primes, a large literature, and numerical data, see \cite{KW1983}, et cetera. 
\section{Representation of the Characteristic Function} \label{S3242.200}
The standard representation of the characteristic function of primitive elements in finite fields below in Lemma \ref{lem3242.900} is divisor dependent.

\begin{lem} \label{lem3242.900}
Let \(G\) be a finite cyclic group of order \(p-1=\# G\), and let \(0\neq u\in G\) be an invertible element of the group. Then
\begin{equation}
\Psi (u)=\frac{\varphi (p-1)}{p-1}\sum _{d \mid p-1} \frac{\mu (d)}{\varphi (d)}\sum _{\ord(\chi ) = d} \chi (u)=
\left \{\begin{array}{ll}
1 & \text{ if } \ord_p (u)=p-1,  \\
0 & \text{ if } \ord_p (u)\neq p-1. \\
\end{array} \right .
\end{equation}
\end{lem}

\begin{proof} A complete proof appears in  Lemma 3.2 in \cite{CN2017}.
\end{proof}

The works in \cite{DH1937}, and \cite{WR2001} attribute the above formula to Vinogradov. The proof and other details on the characteristic function are given in \cite[p. 863]{ES1957}, \cite[p.\ 258]{LN1997}, \cite[p.\ 18]{MP2007}. The characteristic function for multiple primitive roots is used in \cite[p.\ 146]{CZ1998} to study consecutive primitive roots. In \cite{DS2012} it is used to study the gap between primitive roots with respect to the Hamming metric. And in \cite{WR2001} it is used to prove the existence of primitive roots in certain small subsets \(A\subset \mathbb{F}_p\). In \cite{DH1937} it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b \in \mathbb{F}_p$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in \cite{PS1995} using this formula.\\

The result in Lemma \ref{lem3242.300} provides a divisor-free of the representation of the characteristic function of primitive elements in finite fields $\F_p$. 

\begin{lem} \label{lem3242.300}
Let \(p\geq 2\) be a prime, and let \(\tau\) be a primitive root mod \(p\). If \(u\in\mathbb{F}_p\) is a nonzero element, then
\begin{equation}\label{eq3242.300}
\Psi (u)=\sum _{\gcd (n,p-1)=1} \frac{1}{p}\sum _{0\leq k\leq p-1} e^{i2 \pi \left (\tau ^n-u\right)k/p}
=\left \{
\begin{array}{ll}
1 & \text{ if } \ord_p (u)=p-1,  \\
0 & \text{ if } \ord_p (u)\neq p-1. \\
\end{array} \right .\nonumber
\end{equation}
\end{lem}

\begin{proof} A complete proof appears in  Lemma 3.3 in \cite{CN2017}.
\end{proof}

\section{Evaluation Of The Main Term} \label{S3262.000}
The precise evaluation of the main term $M(x)$ occurring in the proof of Theorem \ref{thm3232.100} is recorded here. The symbol $\li(x)$ denotes the logarithm integral.

\begin{lem}  \label{lem3262.300}  Let \(x\geq 1\) be a large number, and let $p$ be a prime. Then,
	\begin{equation} \label{eq3262.010}
	\sum_{p\leq x} \frac{1}{p}\sum_{\gcd(n,p-1)=1} 1=a_1\li(x)+O\left(\frac{x}{\log
		^bx}\right) ,
	\end{equation} 
	where $b>1$ is a constant.
\end{lem}

\begin{proof} A complete proof appears in Lemma 5.1 in \cite{CN2017}. 
\end{proof}
The average density of primitive roots modulo $p$, known as Artin constant, is given by 
\begin{equation} \label{eq3262.020}
a_1=\prod_{p \geq 2 } \left(1-\frac{1}{p(p-1)}\right)= 0.37395581361920228805\ldots, 
\end{equation}  
see \cite{SP1969}, and \cite{WJ1961}.

\section{Estimate For The Error Term} \label{S3282.000}
The upper bounds for the error term $E(x)$ in the proof of Theorem \ref{thm3232.100} is recorded here. 

\begin{lem} \label{lem3282.000}  Let \(p\geq 2\) be a large prime, and let \(\tau\) be a primitive root mod \(p\). If the element \(u\ne 0\) is not a primitive root, then, 
	\begin{equation} \label{eq3282.020}
	\sum_{x \leq p\leq 2x}
	\frac{1}{p}\sum_{\gcd(n,p-1)=1,} \sum_{ 0<k\leq p-1} e^{i2 \pi \left (\tau ^n-u\right)k/p}\ll \frac{x^{1-\varepsilon}}{\log x}\nonumber
	\end{equation} 
	for all sufficiently large numbers $x\geq 1$ and an arbitrarily small number \(\varepsilon <1/16\).
\end{lem}

\begin{proof}  A complete proof appears in Lemma 6.1 in \cite{CN2017}.
\end{proof}

\section{Main Result} \label{S3292.000}
Given a large number $x\geq1$, and a fixed integer $q \ne \pm 1, n^2$, the precise primes counting functions are defined by 
\begin{equation} \label{eq3292.900}
\pi(x)=\# \{ p \leq x: p \text{ is prime} \},
\end{equation}
and
\begin{equation} \label{eq3292.910}
\pi_{q}(x)=\# \{ p \leq x: \ord_p(q)=p-1 \}.
\end{equation}
The density of the subset of primes $\mathcal{D}=\{ p \leq x: \ord_p(q)=p-1 \}$ with a fixed primitive root $q\ne \pm 1,n^2$, is defined by the limit
\begin{equation} \label{eq3292.920}
\delta(q)=\lim_{x \to \infty} \frac{\pi_{q}(x)}{\pi(x)}=c(q)a_1.
\end{equation}
The constant $a_1>0$ is defined in \eqref{eq3262.020}, and the correction factor $c(q)\geq 0$ was computed in \cite[p.\ 220]{HC1967}.

\begin{proof} {\bf (Theorem \ref{thm3232.100})}. Let \(x\geq x_0\) be a large number, and let $z=(\log x)^c$, where $c\geq 0 $ is constant. Suppose that the integer $q \ne \pm 1,n^2$ is not a primitive root for all primes short interval \(p \in [z,2z]\). Summing of the characteristic function of primitive roots in the prime finite field $\mathbb{F}_p$ over the short interval \([z,2z]\) returns the nonexistence equation
\begin{equation} \label{eq3292.940}
	0=\sum_{z \leq p\leq 2z} \Psi (q).
\end{equation}
Replacing the characteristic function, Lemma \ref{lem3242.300}, and expanding the nonexistence equation \eqref{eq3292.940} yield
\begin{eqnarray} \label{eq3292.950}
0&=&\sum_{z \leq p\leq 2z} \Psi (q) \\ 
	&=&\sum_{z \leq p\leq 2z} \left (\frac{1}{p}\sum_{\gcd(n,p-1)=1,} \sum_{ 0\leq k\leq p-1} e^{i2 \pi \left (\tau ^n-q\right)k/p} \right ) \nonumber\\
	&=&\delta(q)\sum_{z \leq p\leq 2z}\frac{1}{p}\sum_{\gcd(n,p-1)=1} 1  +\sum_{z \leq p\leq 2z}
	\frac{1}{p}\sum_{\gcd(n,p-1)=1,} \sum_{ 0<k\leq p-1}e^{i2 \pi \left (\tau ^n-q\right)k/p}\nonumber\\
	&=&M(z) + E(z)\nonumber,
	\end{eqnarray} 
	where $\delta(q) \geq 0$ is a constant depending on the integer $q\geq 2$. \\
	
	The main term $M(z)$ is determined by a finite sum over the trivial additive character \(\psi(t) =e^{i 2\pi  kt/p}=1 \) for $k\ne0$, and the error term $E(z)$ is determined by a finite sum over the nontrivial additive characters \(\psi(t) =e^{i 2\pi k t/p}\ne1 \) for $k\ne0$.\\
	
Applying Lemma \ref{lem3262.300} to the main term, and Lemma \ref{lem3282.000} to the error term yield
	\begin{eqnarray} \label{eq3292.960}
	0&=&\sum_{z \leq p\leq 2z} \Psi (q) \nonumber \\
	&=&M(z) + E(z) \nonumber\\
	&=&\delta(q)\left (\li(2z)-\li(z) \right )+O\left(\frac{z}{\log^bz}\right)+O\left( \frac{z^{1-\varepsilon}}{\log z} \right) \nonumber\\
	&=&\delta(q)\left (\li(2z)-\li(z)\right )+O\left( \frac{z }{\log^b} \right)  ,
	\end{eqnarray} 
	where $\delta(q)\geq 0$ is defined in \eqref{eq3292.920}. \\ 
	
Since the density $\delta(q)> 0$ for $q \ne \pm 1, n^2$, see \cite[p.\ 220]{HC1967}, and the difference of logarithm integrals
\begin{equation}
 \li(2z)-\li(z)\gg \frac{z}{\log z}>0 ,  
\end{equation}
the expression
\begin{eqnarray} \label{eq3292.970}
	0&=&\sum_{z \leq p\leq 2z} \Psi (q)  \\
	&=&\delta(q)\left (\li(2z)-\li(z)\right )+O\left( \frac{z }{\log^b} \right)  \nonumber\\
	&\gg& \frac{z}{\log z}\nonumber\\
    &>&0,\nonumber
	\end{eqnarray} 
is false for all sufficiently large numbers $z=2(\log x)^c$, and contradicts the hypothesis \eqref{eq3292.940}. Ergo, the short interval $[z,2z]$ contains a small primes $p\leq 2z=2(\log x)^c$ such that the $q$ is a fixed primitive root. 
\end{proof}


\section{Open Problems}\label{EXE3535}

\begin{exe}\label{exe3535.0123} {\normalfont Develop a divisor-free primitive root test. The standard primitive root test, see Lemma \ref{lem3242.500}, is totally dependent on the divisors of $p-1$.
}
\end{exe}


\currfilename.\\

\end{document}